A ball is thrown vertically upwards at 5 m/s from the roof of a 100 m building. A second ball is thrown downwards from the same point 2 seconds later at 20 m/s. a) Where and when will the balls meet? b) What are their velocities at that instant?
1
Expert's answer
2021-10-24T18:17:50-0400

"t_1=\frac vg=0.5~s,"

"t=t_2-t_1=1.5~s,"

"s=\frac{v^2}{2g}=1.25~m,"

a)

"s'=s+h-\frac{gt'^2}2,"

"s'=h-v'(t'-t)-\frac{g(t'-t)^2}2,\implies"

"t'=\frac{v't-\frac{gt^2}2+s}{v'-gt}=4~s,"

"s'=21.25~m,"

b)

"v_{b1}=gt'=40~\frac ms,"

"v_{b2}=v'+g(t'-t)=65~\frac ms."